Lesson plan

1.Write down how you would explain the x and y to learners.

2.the relationship between geometrical and number sentences.
Explain how the relationship between the two types of patterns was made in the lessons. Suggest any improvement that you would implement to make the relationship clearer.
3.representational forms
-list the equivalent representational forms for representing pattern as they were used in the two lessons.
4.problem solving
-critically evaluate the problem solving examples in the last part of the second lesson.
1
Expert's answer
2020-07-23T17:53:41-0400

[1 .] They are variables, unknown numbers. Often they are used to refer to axes on the Cartesian plane with x being horizontal and y being vertical.

A “function” is formally defined in mathematics as “a set of ordered pairs of numbers (xy)

[2 .] Numeric sequences we only add and subtract to get the next answer.

Geometric sequences we multiply and divide to get the next answer.

we mostly use relationships where the inputs and outputs are numbers, and where a single input number gives a single output number.

The input and output are often written as an ordered pair.

[3 .] List of represent patterns :

(A). Arithmetic Sequence.

(B). Geometric Sequence.

(C). Square Numbers.

(D). Cube Numbers.

(E). Triangular Numbers.

(F). Fibonacci Numbers.

[4 .] problem solving :

Example of Geometrical Sequence Pattern is : 4 ,44 , 484 , 5324

4 x 11 = 44

44 x 11 = 484

484 x 11 = 5324

The constant ratio between the numbers are x (11) which means that this is a geometric sequence.

Example of Numberic Sequence Pattern : 1 , 3 , 5 , 7

1+2 =3

3+2= 5

5+2 =7

The constant difference for this sequence is +2 which means it is a numeric sequence.